### v1

# Advanced Encryption Standard AI artificial intelligence

AES
Advanced Encryption Standard
AI
artificial intelligence
ALU
arithmetic logic unit
AOI
and-or-invert
API
application programming interface
ARX
add-rotate-XOR
ASIC
application-specific integrated circuit
ASIP
application specific instruction-set processor
AS
active serial
BFS
breadth-first search
BGA
ball grid array
BRAM
block-RAM
CFG
control flow graph
CLB
configurable logic block
CPU
central processing unit
DC
direct current
DDR
double data rate
DFA
differential frequency analysis
DFT
discrete Fourier transform
DSO
digital storage oscilloscope
DSP
digital signal processing
DUT
device under test
EDA
electronic design automation
EEPROM
electrically erasable programmable ROM
EMA
electro-magnetic emanation
EM
electro-magnetic
FFT
fast Fourier transformation
FF
flip-flop
FIB
focused ion beam
FIFO
first in first out
FIR
finite impulse response
FPGA
field-programmable gate array
FSM
finite state machine
GNN
graph neural network
GPIO
general purpose input/output
GUI
graphical user interface
HDL
hardware description language
HD
Hamming distance
HF
high frequency
HRE
hardware reverse engineering
HSM
hardware security module
HW
Hamming weight
I2C
Inter-Integrated Circuit
I2S
Inter-IC Sound
IC
integrated circuit
ICAP
Internal Configuration Access Port
IIR
infinite impulse response
I/O
input/output
IOB
Input Output Block
IoT
Internet of things
IP
intellectual property
IPSW
iPod Software
ISA
instruction set architecture
IV
initialization vector
JTAG
Joint Test Action Group
LFSR
linear feedback shift register
LRA
linear resonant actuator
LSB
least significant bit
LUT
look-up table
MAC
Multiply-Accumulate
MIPS
Microprocessor without Interlocked Pipeline Stages
ML
machine learning
MMIO
memory mapped input/output
MUX
multiplexer
MSB
most significant bit
NASA
National Aeronautics and Space Administration
NDA
non-disclosure agreement
NMI
normalized mutual information
NSA
National Security Agency
NVM
non-volatile memory
OFB
Output Feedback Mode
OISC
One Instruction Set Computer
ORAM
Oblivious Random Access Memory
OS
Operating System
OSI
Open Systems Interconnection
PAR
place-and-route
PCB
printed circuit board
PC
personal computer
PDK
process design kit
PI
proportional-integral
PID
proportional-integral-derivative
PLB
Programmable Logic Block
PS
passive serial
PUF
physical unclonable function
RAM
random-access memory
RE
reverse engineering
RISC
reduced instruction set computer
R&D
research and development
RNG
random number generator
ROM
read-only memory
RTL
register transfer level
SAT
Boolean satisfiability problem
SCA
side-channel analysis
SCC
strongly connected component
SEM
scanning electron microscope
SE
symbolic execution
SHA
Secure Hash Algorithm
SMT
satisfiability modulo theories
SNR
signal-to-noise ratio
SoC
system-on-chip
SPA
simple power analysis
SPI
Serial Peripheral Interface
SRAM
static random access memory
SRE
software reverse engineering
STG
state transition graph
UART
Universal Asynchronous Receiver Transmitter
UHF
ultra-high frequency
USB
Universal Serial Bus
VHDL
Very High Speed Integrated Circuit Hardware Description Language
[…]
Stealing Maggie’s Secrets—On the Challenges of IP Theft
Through FPGA Reverse Engineering
Simon Klix1[Uncaptioned image], Nils Albartus1[Uncaptioned image], Julian Speith1[Uncaptioned image], Paul Staat1[Uncaptioned image], Alice Verstege2[Uncaptioned image], Annika Wilde2[Uncaptioned image],
Daniel Lammers2[Uncaptioned image], Jörn Langheinrich1[Uncaptioned image], Christian Kison3[Uncaptioned image], Sebastian Sester3[Uncaptioned image],
Daniel Holcomb4[Uncaptioned image], Christof Paar12[Uncaptioned image]
1Max Planck Institute for Security and Privacy, Bochum, Germany 2Ruhr University Bochum, Horst Görtz Institute for IT Security, Bochum, Germany 3Bundeskriminalamt, Wiesbaden, Germany 4UMass Amherst, MA, USA
Abstract
Intellectual Property (IP) theft is a cause for major financial and reputational damage, reportedly in the range of hundreds of billions of dollars annually in the U.S. alone. Field Programmable Gate Arrays (FPGAs) are particularly exposed to IP theft, because their configuration file contains the IP in a proprietary format that can be mapped to a gate-level netlist with moderate effort. Despite this threat, the scientific understanding of this issue lacks behind reality, thereby preventing an in-depth assessment of IP theft from FPGAs in academia. We address this discrepancy through a real-world case study on a Lattice iCE40 FPGA found inside iPhone 7. Apple refers to this FPGA as Maggie. By reverse engineering the proprietary signal-processing algorithm implemented on Maggie, we generate novel insights into the actual efforts required to commit FPGA IP theft and the challenges an attacker faces on the way. Informed by our case study, we then introduce generalized netlist reverse engineering techniques that drastically reduce the required manual effort and are applicable across a diverse spectrum of FPGA implementations and architectures. We evaluate these techniques on seven benchmarks that are representative for different FPGA applications and have been synthesized for Xilinx and Lattice FPGAs. Finally, we provide a comprehensive open-source tool-suite of netlist reverse engineering techniques to foster future research, enable the community to perform realistic threat assessments, and facilitate the evaluation of novel countermeasures.
Index Terms:
IP theft, reverse engineering, bitstream, gate-level netlist, FPGA, iPhone
1 Introduction
Many companies invest heavily into creating intellectual property (IP), which often forms the backbone of their economic value and competitive advantage. For instance, Apple alone spent over \$26 billion on research and development (R&D) in 2022 [1]. Not surprisingly, valuable IP can nowadays often be found within electronics and integrated circuits in particular. Here, it can take the form of proprietary algorithms for domains like digital signal processing (DSP), CPUs, or security functions. As a flip side of this development, IP theft, e.g., by competitors or hostile nation states, has become a major issue. In the U.S. alone, IP theft causes hundreds of billions of dollars in damages every year, with the semiconductor market being of particular concern [2]. In this paper, we look at an important aspect of this issue, namely assessing the security of IP in field-programmable gate arrays, which are, as we argue, particularly vulnerable against IP theft.
In many highly specialized domains—such as spaceflight, military communications, network routers, or medical devices—FPGAs are a vital component, often incorporating proprietary IP. FPGAs are (re-)configurable logic devices that are programmed using a bitstream with a proprietary file format. This bitstream is generated by electronic design automation (EDA) tools that are provided by the FPGA vendors themselves. These tools map a high-level design, written in a hardware description language (HDL) such as Verilog, to a bitstream during synthesis. At the end of this process, the implemented IP is fully encapsulated in the bitstream file. This file is then stored externally to the FPGA, but residing on the target device.
To better assess the threat of, and defend against, real-world IP theft from FPGA implementations, it is crucial to improve our understanding of the associated challenges and required efforts. In most cases, hardware IP theft entails reverse engineering the design in order to recover the implemented algorithms and their instantiation.
Modern nanometer-scale ICs come with a natural defense against reverse engineering due to the high cost and labor-intensive nature of the netlist extraction process [3, 4, 5, 6, 7]—a financial barrier that deters all but the most resourceful entities, such as nation-states. However, the bar for FPGA reverse engineering is much lower. To this end, an attacker would need to gain access to the bitstream and map it to a gate-level netlist, a process that is already well understood in research despite the proprietary nature of the bitstream format [8, 9, 10, 11, 12, 13, 14, 15]. FPGA vendors commonly provide cryptographic bitstream protections as a safeguard for valuable IP. Nonetheless, these measures have repeatedly been demonstrated to be susceptible to side-channel attacks [16, 17, 18, 19, 20] or protocol flaws [21, 22]. Interestingly, many low-cost FPGAs do not even offer such protections. Even if these protective measures are available and would withstand such attacks, they must be proactively enabled by the FPGA developer. Hence, we assume access to the plaintext bitstream, and thus the gate-level netlist, is generally achievable in practice.
In this situation, a reverse engineer is typically faced with a gate-level netlist that lacks any hierarchy, module boundaries, and word-level information such as data types, see Section 2.3. Many netlist reverse engineering techniques have tried to address these challenges in recent years [23], but they often deal only with isolated sub-problems such as register recovery, lack reference implementations, and use unrealistic and/or outdated benchmarks, see Section 5.3. In particular, an end-to-end analysis of a real-world FPGA design remains largely unexplored in literature. As a consequence, the actual threat potential of FPGA IP theft as well as the unique challenges it entails remain unknown. Thus, the motivation of the paper at hand is to investigate the reverse engineering of a complex FPGA netlist that is recovered from a real-world device in a black-box setting. The goal is to obtain and analyze the algorithmic description of the design. A better understanding of this issue can aid in the design of sound defenses and inform risk analysis of industry as well as governments. We note that FPGA reverse engineering also paves the way for attacks other than IP theft, e.g., retrieving hard-coded cryptographic keys or inserting hardware Trojans.
Research Questions and Contributions
For a realistic assessment of the challenges of real-world FPGA reverse engineering in the context of IP theft, we first set out to answer the following research question:
RQ1 What challenges and efforts are entailed with real-world IP theft through FPGA reverse engineering in a black-box setting?
To answer RQ1, we demonstrate the steps and efforts required to extract IP from an FPGA implementation through a case study on a Lattice iCE40 FPGA found within the iPhone 7. Apple refers to this FPGA as Maggie in the iPhone firmware. To this end, we combine existing techniques and propose new approaches to recover an algorithmic description of the target FPGA implementation, effectively extracting the IP controlling the iPhone’s Taptic Engine, see Section 3. In addition to evaluating existing techniques and proposing novel approaches to FPGA reverse engineering, we provide unique insights that foster developing better defenses against IP theft. Based on our findings, a second research question arises naturally: RQ2 To what extent can FPGA reverse engineering be generalized across architectures and implementations?
In response to RQ2, we generalize the methods developed for our case study. This yields automated methods that are applicable across FPGA architectures and implementations. In particular, we target Xilinx 7-Series and Lattice iCE40 FPGA families, see Section 4. We evaluate the presented techniques on a diverse set of benchmark netlists and discuss differences in effectiveness. In line with general efforts of the security community to provide open-source implementations, thereby removing financial barriers and enabling third-party verification, we will share our benchmarks and implementations as plugins to HAL [24] upon publication of this work. Thereby, we hope to encourage more insight into FPGA reverse engineering in follow-up work by academia and industry, and foster the development of countermeasures, see Section 5.
Responsible Disclosure
Despite not revealing any security vulnerabilities, we informed Apple and Lattice Semiconductor about our findings prior to publication.
2 Attacker Model & Challenges
We now present our attacker model, introduce relevant terminology, and discuss the challenges of netlist reverse engineering encountered in our case study.
2.1 Attacker Model
Our attacker model revolves around IP theft, which is a common threat considered in hardware security research [25, 26]. An attacker obtains a device featuring an FPGA and aims to recover the contained IP by extracting an algorithmic model of the design implemented on the FPGA by means of reverse engineering. We assume the attacker has gained access to the bitstream configuring the FPGA by either extracting it from the device or from firmware. Given that the attacker possesses the device, they can execute the FPGA implementation and modify the hardware as needed.
As is typical in reverse engineering, the attacker may have already gathered public information on the target device, such as documentation or patents. This leaves them with knowledge on the system’s connectivity and interaction with the FPGA, along with a general understanding of its purpose, but lacking detailed knowledge of the implementation and the intricacies of the algorithm. Ultimately, the attacker seeks to reconstruct a detailed algorithmic representation that allows him to extract, understand, and modify the IP contained within it. Thus, our attacker model explicitly goes beyond simply copying an FPGA bitstream or netlist.
2.2 Netlist Reverse Engineering Terminology
A gate-level netlist is a digital circuit representation comprising combinational and sequential gates or standard cells as well as their interconnections, also known as nets. In reverse engineering, we frequently encounter flattened netlists lacking hierarchy or module boundaries. Furthermore, we differentiate between data path and control path. The control path steers the data path in that it decides which operations are being performed on the data passing through by enabling or disabling certain parts of the circuit.
Reverse engineering gate-level netlists involves static and dynamic techniques. Static methods analyze the netlist graph and can take either structural or functional information into account. Dynamic techniques leverage data gathered during execution to reverse engineer functionality.
2.3 Challenges of Netlist Reverse Engineering
Committing IP theft on an FPGA requires analyzing the gate-level netlist extracted from the bitstream, posing unique challenges. These challenges arise from the need to reconstruct information removed during synthesis. The challenges we faced in our case study include:
  1. C1
    Loss of Hierarchy. HDL designs feature modules that encapsulate functionality and introduce hierarchy. This information is removed during synthesis. The resulting netlist lacks any module boundaries or hierarchy.
  2. C2
    Loss of Data Types. An HDL description makes use of multi-bit data types like integers on which, e.g., arithmetic operations are performed. In contrast, gate-level netlists only contain gates and nets operating on bit level. Not only is the information of which gates belong to the same word-level operation lost during synthesis, but also the order of bits within word-level structures. Consequently, high-level data types no longer exist in the netlist which hampers analysis relying on, e.g., integer values.
  3. C3
    Synthesizer Optimizations. Synthesizer optimizations can result in similar operations on HDL level looking vastly different in a gate-level netlist. Combinational logic may be merged across module boundaries and gates or even entire sub-circuits may be duplicated to facilitate a high fan-out. Such optimizations always depend on the available standard cells, timing and area constraints, routing requirements, and varying optimization strategies.
  4. C4
    Missing Control Separation. The control path is often implemented using finite state machines. In an HDL design, FSMs can be distinguished from one another and the data path. In netlists, FSMs are often merged with one another or even the data path such that boundaries can no longer be determined. This causes a state explosion when analyzing affected FSM state graphs, sometimes raising complexity beyond what is feasible.
  5. C5
    Data Dependency. The behavior of the control path—and in extension the data path—depended on external data and instructions fed to the FPGA as input. Hence, analyzing such a netlist without input data is challenging, especially if the control path cannot be properly dissected.
  6. C6
    Dynamic Behavior. Sequential gates add another dimension to netlist reverse engineering in that they require the reverse engineer to consider past states for the analysis of current behavior. This blows up complexity due to the vast number of possible netlist states to consider.
  7. C7
    Semantic Analysis. Extracting a word-level algorithmic representation leaves a reverse engineer with the task of assigning meaning and symbols to different values of the computation. To gain a high-level understanding, including the underlying rationale and specific design decisions, the functionality needs to be dissected rigorously from the system-level down to individual building blocks, calling for domain-specific expertise.
3 Case Study on iPhone 7
To answer RQ1 and investigate the effort required to commit IP theft from a real-world FPGA implementation, we conducted a case study on an FPGA found in iPhone 7.
Target Device
The iPhone 7 and iPhone 7 Plus were released in September 2016. The smartphones feature a novel capacitive home button with haptic feedback being generated by the Taptic Engine first introduced in Apple Watch and iPhone 6s. Notably, both iPhone 7 and iPhone 7 Plus contain a Lattice iCE5LP4K-SWG36 FPGA. Apple refers to this chip as Maggie throughout firmware and leaked printed circuit board (PCB) schematics. There are many speculations revolving around the purpose of Maggie [27, 28, 29].
Figure 1: Maggie as part of the Taptic Engine controller.
Setting
From analyzing leaked PCB schematics, the iPhone firmware, and patents granted to Apple [30, 31], we concluded that Maggie controls the Taptic Engine, and determined its I/O connectivity and communication interfaces, see Figure 1. This alone does not enable effective IP theft, as we still lack implementation details on the instantiated signal processing algorithms and their parameters. Assuming the role of the attacker, we therefore need to investigate the circuit implemented on Maggie to fully recover Apple’s Taptic Engine IP. To this end, we split the iPhone FPGA reverse engineering process into four steps, see Figure 2.
Figure 2: Overview of our case study on iPhone 7.
3.1 Step 1: Netlist Recovery
At first, we were faced with the iPhone 7 containing the bitstream featuring the target IP and our first goal was to extract the gate-level netlist for further analysis.
Bitstream Acquisition
SRAM-based FPGAs such as Lattice iCE40 devices store the bitstream externally and are configured during startup. Hence, an attacker can read the bitstream from external memory, intercept it during transmission, or recover it from device firmware. The Maggie bitstream is shipped with the openly available iPhone firmware, hence we extracted it from the file system. The original bitstream was published in September 2016 and was updated with iOS 10.2 in December 2016. As it remains unchanged ever since, our analysis deals with this latest version.
Bitstream Format Reverse Engineering
The bitstream format is typically a secret well-kept by FPGA vendors, although everyone can generate bitstreams using respective EDA tools. Project IceStorm [32] entails a comprehensive documentation database of the Lattice iCE40 bitstream format generated by fuzzing the Lattice iCEcube2 EDA tool. Project IceStorm is part of F4PGA [33], an open-source EDA tool-flow developed as an alternative to proprietary vendor software. As part of our efforts, we extended Project IceStorm by increasing its fuzzing coverage and adding support for the pinout of Maggie’s SWG36 chip package.
Bitstream Conversion
Given a database detailing the mapping from bits to netlist elements, an arbitrary bitstream can be converted back to a gate-level netlist in full automation. In light of our case study, we developed a custom bitstream conversion tool that translates any Lattice iCE40 bitstream into a gate-level netlist. Our tool depends on the Project IceStorm database to be complete and correct. For Maggie, we observed 36 nets with missing sources or destinations, hinting at an incomplete bitstream database. Later on, we manually repaired these nets using information on surrounding circuitry revealed once our analysis progressed.
The extracted netlist contains 5046 gates comprising 3241 look-up tables, 422 carry gates, 1342 flip-flops, 18 block-s, and 4 DSP units.
3.2 Step 2: Word-Level Reconstruction
After extracting the gate-level netlist, we now aim to recover word-level structures from the unstructured sea-of-gates using the netlist reverse engineering framework HAL [24]. To this end, we were confronted with the challenges outlined in Section 2.3.
Netlist Pre-Processing
To eliminate synthesis byproducts hindering analysis and enhance readability for human reverse engineers, we initially pre-processed the netlist.
(a) Replace LUTs with primitive combinational gates.
(b) Decompose LUTs into a MUX and primitive gates.
Figure 3: Netlist Preparation Overview.
For more efficient manual analysis and considering that some LUTs merely implement basic functions like AND or XOR, we substituted such LUTs with primitive gates, see 2(a). We also removed buffer LUTs as they add no functionality but sometimes hinder structural analysis. Furthermore, we removed gates and sub-circuits that compute the same function on identical inputs. Such duplicates are introduced by the synthesizer for optimization purposes, yet they can obfuscate shared control signals or data inputs.
MUXes in particular provide structure to the data path. On FPGAs, however, their functionality is implemented as part of more complex Boolean functions within LUTs. Hence, decomposing complex LUTs into small building blocks such as MUXes can aid manual and automated reverse engineering, e.g., by supporting structural analysis. To recover a nested MUX and better understand the data path (see Figure 5), we analyze the LUT’s Boolean function and search for inputs that behave like a select. When discovering a MUX in a LUT, we replace the LUT with a MUX and reconstruct the surrounding logic from primitive gates.
In total, we replaced 549 LUTs with primitive gates, removed 594 buffers, extracted 1619 MUXes and 1161 primitive gates, removed 353 duplicate gates, and removed 51 additional gates using other simplifications.
Control Logic
Analyzing Maggie’s control path was required for comprehending how the circuit operates. Additionally, certain techniques operating on the data path require knowledge on which gates and nets belong to the control path. Unfortunately, the absence of FPGA-compatible reference implementations [34, 35, 36], the absence of public implementations altogether [37, 38, 39, 40], or general assumptions on strongly connected components that do not always hold for our case study [39, 38], prevented us from applying existing techniques from literature. In particular, the NETA toolset [41] is only available as binaries that cannot parse FPGA netlists. Moreover, a common assumption is that the state registers and transition logic of an FSMs form a SCC [39, 38]. For Maggie, we found 55 SCCs with all but seven comprising three gates or less. Vast parts of the netlist form a single SCC of more than 5000 gates. Therefore, SCCs are not a useful metric in our case study.
Given the intricacies of this issue, we ourselves could not find a workable solution within the scope of this work and hence decided to manually investigate the control path of Maggie instead. Here, we faced significant challenges due to the FSMs being interwoven to an extent that often prevented even manual separation, as discussed in [40]. While FSM separation was not always feasible, we still classified FFs as part of the control path by checking whether they feed into control pins such as MUX select or FF enable. After assigning a FF to the control path, we also examined its predecessors and successors. These structural characteristics are merely clues that a gate might belong to the control path. They often still required manual inspection and human intuition, hence preventing full automation.
Communication Interfaces
Investigating Maggie’s SPI, UART, and I2S interface implementations (see Figure 1) supports word-level structure recovery as these interfaces often comprise the first or last register of a data path. Furthermore, we must reverse engineer these interfaces to understand the parallel and serial conversion of incoming and outgoing data, which is precondition for recovering bit orders and data types. Through manual reverse engineering, we identified 615 gates belonging to five interfaces.
Arithmetic Structures
In the iPhone netlist, we observed the data path comprising mostly combinational logic implemented as LUTs and carry gates as well as sequential logic implemented using FFs. Some data-path structures are implemented in a consistent, recurring way, e.g., adders, counters, and selected comparators. Lattice iCE40 FPGAs provide carry gates that are commonly connected to form carry chains, often used in combination with LUTs to implement arithmetic operations, see Figure 4. A reverse engineer can leverage these recurring pattern to identify word-level structures like adders or counters in a gate-level netlist.
Figure 4: Structure of a $3$-bit counter with reset constructed from a carry chain on a Lattice iCE40 FPGA.
To recover word-level arithmetic operations, we first locate carry chains in the netlist by scanning for connected carry gates. We then assemble arithmetic structure candidates by adding surrounding logic to the carry chains. Since a structure may be influenced by control inputs, we construct structural candidate variants to account for different control behavior. Operand membership as well as the order of input and output signals are derived from the structure and connectivity of a candidate’s gates. For example, the order of the input signals is reconstructed by identifying the carry gate of the chain that a signal is connected to. Finally, we check each candidate against predefined satisfiability modulo theories (SMT) models, e.g., $A\pm B$ for adders or $A\pm n$ for counters incrementing by a constant $n$.
Sequential Data Path
The synthesizer realizes word-level registers from the HDL design as individual FFs that lack any indication of which register they belong to. To this end, DANA [42] can be used to recover these word-level registers from a gate-level netlist using structural metrics such as shared predecessors and successors as well as common control signals. However, as DANA is optimized for data-path analysis, control FFs may not be recognized as such and instead be wrongfully merged into data-path registers. This impairs the detection of other registers due to DANA’s iterative nature. Hence, we extended DANA to include known registers identified by other analysis techniques for finding, e.g., communication interfaces, FSMs states, and counters. DANA does not alter these registers, but utilizes them to identify registers further up or down the data path.
Combinational Data Path
In the data path, MUXes can be used to switch between word-level data sources, such as global inputs, memory, or arithmetic structures. We implemented a method to group MUXes that have been decomposed from LUTs as part of our pre-processing, see Figure 3. A MUX selects an input to be forwarded to the output based on a select input S. Multiple MUXes sharing a common select typically form a word-level MUX structure. Detecting such structures allows for the separation of word-level data paths, see Figure 5. However, grouping by select can yield oversized structures as the same select signal may influence two or more distinct data paths. Hence, we split these groups according to preceding or successive word-level structures. For example, if a 32-bit MUX is succeeded by two 16-bit registers, the MUX is split into two 16-bit MUXes. As other preceding or successive MUXes themselves are considered for splitting, the process is repeated iteratively until no further changes can be observed.
Figure 5: The recovery of word-level MUXes allows for the analysis of separate independent data paths.
Bit Order
While the order of signals (i.e., the bit order) belonging to a word is inherently known in the HDL design, it is removed during synthesis. Hence, for a reverse engineer it is not clear which bits of, e.g., a register are the LSB or MSB. Recovering the bit order of word-level structures aids manual investigation of the design and is vital for the reconstruction of data types used during simulation, see Section 3.3. For adders, counters, BRAMs, and DSPs, the bit order can be inferred from their topology and function. Implementations of other structures, e.g., word-level registers and MUXes, do not exhibit an obvious bit order.
[…]
We propagated the bit order from structures with known bit order along the data path to structures with unknown bit order. For every unordered pin group, our algorithm ($i$) derives bit-order candidates from predecessors and successors and ($ii$) tries to establish consensus between them. Candidates that have an inherent conflict are discarded, e.g., when different pins of a pin group fetched the same index from the same origin. We omit control logic during analysis, as it does not feature a natural bit order and degrades results.
Outcome
Structure #Recovered #Gates Automated?
Arithmetic
35 692
Register
118 1342
Multiplexers
160 1217
Control Logic
- 631
TABLE I: Results of our word-level reconstruction
In addition to the presented methods, we manually grouped 688 gates into data-path modules that did not feature MUXes. In total, we assigned 4570 out of 5496 gates (83%) to structures using manual and automated techniques. Furthermore, we assigned a bit order to 605 (71%) out of 843 pin groups from word-level structures that we did not previously classify as control logic. While we could not verify against a ground truth because of the black-box nature of our case study, these results enabled correctly extracting the implemented algorithm, thereby underlining their value.
3.3 Step 3: Algorithmic Recovery
Having recovered word-level structures in the previous step, we now extract an algorithmic description of what is implemented by the iPhone netlist. This entails further analyzing the control path, as we could not draw boundaries between individual control structures so far.
Virtual Probing
Instead of further attempting to unravel the semantics of the interwoven control logic, we decided to use dynamic reverse engineering techniques to observe the actual behavior of the circuit at runtime. However, such dynamic analysis often requires expensive equipment and yields only incomplete insights into the execution state [43, 20]. Hence, we leveraged a more workable approach: netlist simulation using I/O signals captured during operation. We further used this approach to ($i$) observe the flow of real-world data through the netlist, assisting in algorithmic recovery, and ($ii$) to verify correctness of the recovered netlist and the later extracted algorithm implemented on Maggie by comparing simulation results against recorded outputs.
[…]
We captured I/O signals by placing a custom breakout board between Maggie and the iPhone PCB, see Figure 7. By feeding the captured I/O signals to a simulator, we could then analyze the state of the circuit at any point in time. This reduced complexity of the circuit’s theoretical state space by focusing only on the states that are actually reached. To this end, we extended the netlist analysis framework HAL [24] with simulation capabilities, see Appendix A. This approach allowed for, e.g., analysis of the data flow over time and correct traversal of states for complex FSMs that otherwise exhausted functional analysis. Given the previously reconstructed word-level structures and their bit orders, we could now follow the flow of multi-bit data through the design and observe which operations were executed at what point in time. This completely alleviated us from the need to statically analyze the control path.
DSP Analysis
By tracing inputs to Maggie through the netlist in simulation, we found that almost all data-path operations are performed by four interconnected DSPs. On Lattice iCE40 FPGAs, DSPs feature four 16-bit data inputs and a 32-bit output [44]. On Maggie, all DSPs are configured as Multiply-Accumulate (MAC) units performing 16-bit multiplication followed by 32-bit accumulation. Most DSP inputs are preceded by LUTs implementing MUXes to allow for dynamically changing data sources, see Figure 8. Each DSP and its preceding MUXes are controlled by their own FSM. Looking at the simulation, we found that every DSP computes a recurring sequence of operations depending on dynamically changing control and data inputs.
Figure 8: Simplified DSP circuit.
For each step of these sequences, we leveraged simulation to (i) analyze control inputs and the internal DSP state to determine the executed arithmetic operation and (ii) identify the data sources which the operands of this operation came from. The operands may have passed through MUXes and can originate from communication interfaces, registers, BRAMs, or other DSPs. We also encountered feedback loops when a DSP output is stored in a register that is again applied as an input to the same DSP later on, see Reg1 of Figure 8. Please note that we ignored the actual operand values, but rather traced their data sources to reconstruct symbolic equations describing the executed operations.
For this, we started at the inputs of the respective DSP. Whenever encountering a combinational structure such as a word-level MUX, we looked at the current value of its select signal in simulation to determine the selected data path. Sequential components add another dimension in that they require knowledge of past circuit states, which blows up complexity of static analysis. However, we could leverage information from simulation to avoid investigating all possible circuit states. For registers, we determined the simulation cycle in which they were last written and continued traversal from there. We proceeded with this manual approach until we reached another DSP, a memory, or a communication interface. By applying this to all four DSPs, we generated equations describing Maggie’s data path. Using these equations, we crafted a Python script that replicates the algorithm implemented by the data path and verified correctness by checking against recorded outputs.
3.4 Step 4: High-Level Sense-Making
Based on the recovered algorithm, we proceed with a semantic analysis that is essential for meaningful interaction with Apple’s IP. Generating a Python script implementing the algorithm proved crucial, since comprehending the algorithm required domain knowledge that was beyond our FPGA reverse engineering team. This way, we could share the Python script containing the reconstructed algorithm with a signal processing expert without requiring them to have any knowledge of the reverse engineering process.
Big Picture
Maggie is a central component of the iPhone 7 haptics subsystem to which Apple refers as Taptic Engine. We found that it controls the excitation of the  linear resonant actuator (LRA) to produce a vibration which is perceived by the user as a tactile stimulus. For instance, the Taptic Engine creates the sensation of a physical button push for the non-mechanical home button of the iPhone 7. To provide such refined tactile feedback, a closed-loop control system is utilized which makes the LRA accurately follow a desired movement. Here, Maggie implements the closed-loop motion controller, see Figure 9 for a simplified block diagram, which conditions the drive signals applied to the actuator (via the motor driver) based on sensor feedback from the actuator itself.
Signal Processing
To generate the LRA control signals, Maggie operates on two feedback signals from the LRA, a reference signal previously loaded into BRAM, and values of intermediate signals at previous sampling instants. From the algorithmic representation, we identified several classical biquad filter stages implemented in a direct form I structure [45]. Upon closer inspection, we found many typical DSP building blocks, e.g., sub-sampling and sample-and-hold blocks for sample rate conversion, bit shifts and truncations for fixed-point arithmetic, saturation stages to prevent overflows, and filter implementations. Furthermore, we recognized even subtle details such as the use of fraction saving [46]. Putting it all together, we found three main processing stages: $(i)$ input signal conditioning and application of calibration, $(ii)$ a state observer to track the actuator dynamics, and $(iii)$ the closed-loop controller to minimize errors between the actual and the desired LRA movement. For details on the recovered DSP operations, see Appendix B.
Figure 9: Overview of Maggie’s signal processing.
4 Deriving Generalized Techniques
Our case study demonstrated that a semi-automated approach for FPGA reverse engineering in the context of IP theft entails considerable effort. In line with RQ2, we now investigate whether (and to what extent) the techniques from our case study can be across different FPGA architectures and implementations. We provide open-source implementations of the tools presented in this section as plugins to HAL [24] upon publication of this paper.
We evaluated the techniques discussed in this section on seven benchmark designs synthesized for Xilinx 7-series and Lattice iCE40 FPGAs, see Table II for details on the benchmarks and Table III for evaluation results. Differences in resource utilization between the Xilinx and Lattice benchmarks, e.g., for fft64, can be the result of the Xilinx synthesizer using DSPs where the Lattice synthesizer resorts to carry chains. Furthermore, Xilinx FPGAs leverage 6-input LUTs while Lattice FPGAs use 4-input ones.
4.1 Netlist Pre-Processing
So far, our pre-processing steps were limited to the 4-input LUTs found on Lattice iCE40 FPGAs. While some techniques such as the detection of duplicate gates could easily be extended to cover arbitrary combinational gates, this is not true for our LUT decomposition. Here, our approach works reasonably well for 4-input LUTs, but does not scale to six inputs. Hence, for generalization we re-synthesize the combinational subgraphs between sequential gates using open-source EDA tools instead of analyzing each LUT’s Boolean function. To this end, we constrain the synthesizer to a custom gate library comprising only a the desired combinational gates such as AND, OR, XOR, and MUX of various sizes. Compared to decomposing the combinational parts of the netlist into a pure and-or-invert (AOI) graph, this approach allows us to delegate multiplexer (MUX) detection to the synthesizer by incentivizing the selection of MUX gates over alternatives by artificially reducing their size and optimizing for area. This way, we can even identify MUXes within complex LUT configurations. By nature, this re-synthesis approach is independent of the underlying FPGA architecture and target implementation.
Evaluation
The pre-processing is a best effort approach and lacks any ground truth to compare against. It is, however, a precondition for subsequent techniques such as word-level MUX detection and arithmetic structure identification. The success of these techniques serves as a proxy to gauge the effectiveness of our pre-processing approach.
4.2 Arithmetic Structures
Figure 10: Module Identification and Classification Overview.
In our case study we identified arithmetic operations by building candidate sets of gates and verifying their function using an SMT solver. However, utilizing an SMT solver to compare against functional models from a library of known operations requires knowledge of the order of input and output signals as well as their operand membership. Furthermore, identifying any existing control signals presents additional challenges. When analyzing Maggie, we inferred operands of arithmetic operations from structural properties such as the pins of carry gates, e.g., operand $A$ of an addition $A+B$ would often be connected to input I0 of a Lattice iCE40 CARRY gate. The bit order of arithmetic operations was deduced from the structure of the underlying carry chains. This approach is similar to an idea proposed by Narayanan et al. [47]. However, such structural properties always depend on the particular implementation, FPGA architecture, surrounding logic, and synthesizer optimizations, hence calling for a more generic solution.
[…]
For both Xilinx and Lattice FPGAs, arithmetic operations are often implemented using carry gates. The characteristics of these carry gates can, however, substantially differ between vendors. Hence, different approaches for candidate detection are required depending on the FPGA architecture.
[…]
A salient advantage of this methodology is its generic nature that relieves the user from the intricate task of manually crafting exact candidates. So far, we implemented this structural approach for both Xilinx and Lattice iCE40 platforms. To integrate other FPGA architecture, the structural characteristics of that architecture must be investigated and techniques to compile structural candidates must be implemented. This can be done in a matter of hours.
[…]
Having identified a set of structural candidates, we attempt to determine their implemented operations using only functional methods to remain architecture agnostic. We first determine bit order, operand memberships, and control signals for each structural candidate. Afterwards, we check against functional models from a library of known arithmetic operations using an SMT solver.
[…]
Evaluation
We verified functionality under at least one control assignment for 96% of the carry chains across our benchmarks, see Table III. Issues only arose when carry chains where strongly interwoven with surrounding control and data-path logic, which sometimes prevented extraction of correct structural candidates. Depending on the benchmark, we automatically verified that 3% to 91% (average: 32%) of all combinational gates implement an arithmetic operation. The variation in coverage is the result of different kinds of benchmark implementations. For signal processing designs like the hilbert_transform, canny_edge_detector, and the fft64, the data path mostly comprises of arithmetic operations. central processing unit (CPU) implementations such as ibex, icicle, and simple_risc_v contain fewer carry chains and more control logic, which explains lower numbers in detected arithmetic operations and overall combinational coverage.
4.3 Data Path
The idea behind the dataflow analysis methodology of DANA [42] is, by nature, independent of the target gate type. In our case study, we leveraged DANA to identify registers and developed a separate, but related, technique to group MUX gates into word-level structures. Similar to DANA, we used control signals as well as predecessor and successor gates to identify word-level MUXes. Accordingly, to generalize our data-path analysis, we extended DANA to take on arbitrary, user-defined gate types, including MUXes, instead of refining our initial approach. In addition, we added the option to provide known word-level structures of any kind to DANA to support word-level recovery of the target gate type. For example, in our case study DANA could only search for registers and only previously identified registers could be provided as known groups. Now, DANA can also reconstruct word-level MUXes while taking previously identified registers and arithmetic operations into account.
Evaluation
To evaluate the success of register recovery, we followed the evaluation proposed for DANA [42]. We generated a ground truth by leveraging symbols left in the gate-level netlists of our benchmark by the synthesizer. Then, we use normalized mutual information (NMI) and purity as metrics to compare our refined approach against the original DANA implementation. To this end, we fed known word-level structures such as BRAMs, DSPs, and arithmetic operations to DANA as additional predefined knowledge to take into account during analysis. This marginally improved NMI and purity for most benchmarks, raising the NMI by up to 0.09 the the purity by up to 0.24, see Table III.
We also employed DANA to recover word-level MUXes. However, by nature, no ground truth for MUXes exists, as they are implicitly created from HDL control flow structures by the synthesizer and then implemented in LUTs alongside other logic. Therefore, it is, at best, ambiguous which HDL constructs would be translated into a MUX and accurately measuring the quality of our results is impossible. Still, we provide an intuition for their quality by investigating the sizes of recovered MUXes. MUXes of sizes that divide the bit-size of the datapath or are multiples of a byte are likely to be correct. To this end, we plotted the distributions of the top 5 MUX sizes for all benchmarks in Appendix C.
4.4 Bit Order
Transferring our bit-order propagation algorithm from the case study to other FPGA architectures was straight-forward, because the algorithm was already generic by nature. It operates only on known pin groups of recovered structures and the netlist graph representation to determine successor and predecessor pin groups, which is applicable independent of the underlying gate types. However, CPUs netlists tended to confuse our algorithm due to the interwoven data path of ALUs. To improve bit-order propagation results, we extended our approach for consensus finding between conflicting bit orders propagated to the same pin group. To this end, we now apply three consensus finding mechanisms that try to recover a consecutive bit order.
Consider module with a group of three input pins $[i_{0},i_{1},i_{2}]$. After propagating pin indices from structures with know bit order to those with unknown bit order, each pin is annotated with multiple index candidates coming from different sources. Here, $[0,1,2]$ is a list of indices gathered from the same source structure, denoting that $i_{0}$ should receive index 00, $i_{1}$ index $1$, and $i_{2}$ index $2$. We use ???? to indicate missing index information for the respective pin and $X$ to show that the respective pin is no longer considered.
Shifted Consensus
If conflicting pin orders are propagated to the pins of a pin group, but these orders are just shifted variants of each other, consensus is found and an order starting at index 0 is annotated. For example, if orders $[1,2,3]$ and $[2,3,4]$ are propagated to a group of three pins, the pins will be assigned the order $[0,1,2]$.
Majority Consensus
If the propagated pin orders of a pin group are in conflict, but a non-conflicting majority vote on each individual pin is feasible, the predominant bit order is annotated. For example, if order $[0,1,2]$ is propagated to a group twice and order $[4,3,7]$ is propagated once, $[0,1,2]$ is annotated by majority decision. Since index propagation is performed for each pin individually, there may not be a predominant index for every pin of a pin group. If the majority vote fails for a single pin of a group, no index is annotated to any pin of that group.
Iterative Majority Consensus
Sometimes, the same pin group index is propagated to multiple pins of the same group. For example, the order $[0,?,2]$ may be propagated to a group twice and order [?,1,1]?11[?,1,1][ ? , 1 , 1 ] once. Here, the annotated indices [?,1,1]?11[?,1,1][ ? , 1 , 1 ] will be ignored in the first iteration since they contain the same index for different pins of the same group, therefore conflicting with itself. Instead, indices 00 and $2$ will be annotated to the first and the last pin by majority vote. In the second iteration, only the pins missing an index are considered. The previously gathered groups of indices $[0,?,2]$ and [?,1,1]?11[?,1,1][ ? , 1 , 1 ] are now transformed into $[X,?,X]$ and $[X,1,X]$, thereby resolving internal conflicts. Now $[X,?,X]$ and $[X,1,X]$ cause the second pin to be annotated with index $1$, finally resulting in $[0,1,2]$.
Evaluation
To evaluate our bit-order propagation, we considered the bit orders given for DSPs and BRAMs and the ones recovered for operands of arithmetic operations to be correct and started propagation from there. A reliable bit-order ground truth—given as labels left by the synthesizer—is only available for registers. For evaluation, we created registers based on the ground truth used to evaluate DANA. Furthermore, we only consider registers containing more than three FFs, as smaller ones often belong to the control path. In Table III, we report the absolute number of pin groups we consider for bit-order propagation, the share of initially ordered pin groups, and the total share of ordered pin groups after propagation. In addition, we report the proportion of ordered pin groups that are correct according to our ground truth. On average, we reconstructed 87% of bit orders across our benchmarks with 99% of them being correct, demonstrating the effectiveness of our method. In line with expectations, we observed that the quality of bit-order propagation results improved with the number of structures featuring an inherently known bit order.
4.5 Guided Symbolic Execution
Manually tracing signals across clock cycles in simulation to identify the sources of DSP inputs in our case study turned out not to be scalable as it is tedious and error prone, emphasizing the need for automation. Hence, we implemented a symbolic execution (SE) approach that is guided by concrete control values obtained from virtual probing or simulation to avoid state explosion. This approach can be helpful beyond DSP analysis to semi-automatically investigate a circuit’s dynamic behavior.
Our approach allows symbolically evaluating any data signal and tracing back to its origin to automatically generate equations describing that signal’s behavior over time. These equations depend only on previously defined endpoints such as global inputs, constants (e.g., from memory), or registers. Thus, guided symbolic execution produces a system-level representation that abstracts away timing behavior and instead focuses on the sequence of computations. Finally, the human reverse engineer can manually simplify and interpret the extracted equations to assign high-level meaning.
Starting at the cycle of interest, our guided SE approach first constructs the Boolean function describing the combinational sub-graph in front of a net. Next, it replaces all previously identified control signals with concrete values from simulation while keeping all other variables symbolic. Each variable left in the Boolean function now corresponds to an output net of a sequential gate. If the sequential gate is not one of our endpoints, it finds the last cycle in which the gate was updated and continues back-tracing from the data inputs of the gate at that cycle. This process is repeated until the resulting Boolean function only contains variables corresponding to output nets of our endpoints.
Evaluation
Applying guided SE is highly specific to the target implementation and the goals of the reverse engineer. Therefore, it is hard to evaluate in terms of effectiveness. However, to demonstrate its power, we applied guided SE on the Maggie netlist to extract equations describing the circuit’s algorithmic behavior during execution, after performing the same task manually. Thereby, we semi-automatically extracted the same Python script we previously assembled by hand, see Section 3.3, this time only requiring limited manual intervention. Instead of manually tracing operands back through cycles of the simulation for weeks, we recovered them symbolically using guided SE. Thereby, we reduced the effort required to perform dynamic analysis on Maggie to a matter of days.
TABLE II: Open-source benchmark resource utilization in the number of gates.
Gate Type ibex hilbert_transformer canny_edge_detector icicle simple_risc_v fft64 sha256
Xilinx Lattice Xilinx Lattice Xilinx Lattice Xilinx Lattice Xilinx Lattice Xilinx Lattice Xilinx Lattice
LUT 3503 6379 304 445 2538 3298 1525 2705 1617 2876 1722 5362 2196 3935
Carry 65 295 68 255 636 2170 65 262 76 274 212 2008 104 346
FF 1937 1963 587 569 3713 4484 1009 983 1003 1235 1988 2556 1830 2378
BRAM 0 16 0 2 24 32 270 20 6 6 3 14 0 0
DSP 1 0 0 0 0 0 0 0 0 0 4 0 0 0
Other 727 78 85 78 72 68 78 59 37 31 160 124 269 90
Total 6233 8731 1044 1349 6983 10052 2947 4029 2739 4422 4089 10064 4399 6749
TABLE III: Results of our evaluation on the open-source benchmarks described in Table II and Maggie.
addition   subtraction   counter   negation   const. mul.   comparator   unknown   total   classified   NMI (original)   NMI (our)   purity (original)   purity (our)   no. of groups   initial ordered   final ordered   correct
Design Vendor Arithmetic DANA (Registers) Bit Order
ibex Xilinx
1
0
5
0
0
0
0
6
0.03
0.98 0.98 0.97 0.97 154 0.19 0.86 1.00
Lattice
4
0
4
0
0
0
1
9
0.04
0.96 0.97 0.97 0.97 164 0.25 0.78 0.98
icicle Xilinx
1
1
5
0
0
0
0
7
0.05
0.91 0.95 0.89 0.96 346 0.79 0.88 0.88
Lattice
2
0
7
0
0
1
0
10
0.11
0.85 0.94 0.90 0.92 110 0.30 0.57 1.00
simple_risc_v Xilinx
3
1
4
0
0
2
0
10
0.12
0.97 0.96 0.98 0.95 117 0.29 0.85 1.00
Lattice
2
1
4
0
0
6
1
14
0.15
0.98 0.98 0.98 0.96 136 0.32 0.83 1.00
canny_edge_detector Xilinx
59
0
1
5
46
3
4
118
0.82
0.81 0.81 0.61 0.64 624 0.46 0.90 0.99
Lattice
68
0
2
5
30
5
8
118
0.60
0.81 0.83 0.68 0.77 813 0.43 0.76 1.00
fft64 Xilinx
25
5
0
2
4
0
4
40
0.23
0.89 0.97 0.73 0.97 349 0.36 0.97 0.97
Lattice
93
4
7
0
4
0
1
109
0.44
0.88 0.94 0.76 0.92 622 0.57 0.87 1.00
hilbert_transformer Xilinx
7
10
0
0
0
0
0
17
0.91
0.85 0.95 0.68 0.90 126 0.43 1.00 1.00
Lattice
7
10
0
0
0
0
0
17
0.89
0.90 0.94 0.74 0.85 126 0.46 1.00 1.00
sha256 Xilinx
10
0
0
0
0
0
3
13
0.14
0.91 0.96 0.75 0.87 120 0.28 0.95 1.00
Lattice
11
0
1
0
0
0
0
12
0.10
0.94 0.97 0.81 0.90 170 0.22 0.98 1.00
maggie Lattice
3
0
25
0
0
12
2
42
0.15
N/A N/A N/A N/A 315 0.47 0.78 N/A
5 Discussion
We begin by revisiting our research questions, followed by a discussion on the threats of FPGA reverse engineering, along with challenges and prospects of potential defenses. Finally, we revisit related work and elaborate on limitations and future research directions.
5.1 Revisiting Research Questions
RQ1
In our case study, we successfully recovered the algorithm controlling the Taptic Engine of the iPhone 7. To answer RQ1 and provide insights into the effort required to commit IP theft, we now share our experiences.
While netlist recovery was solved following a known path with suitable open-source tooling, netlist analysis presented challenges that took months to complete. In particular, the lack of hierarchy (C1), loss of data types (C2), and synthesizer optimizations (C3) initially hampered structural analysis, but could be dealt with once patterns were identified. This was achieved through an interplay of manual and automated reverse engineering by first searching for repeating structures and then systematically automating detection thereof. While we automated substantial parts of the word-level reconstruction, manual inspection and correction were inevitably necessary to interpret results and address errors.
Classification of control logic (C4) had to be mostly done by hand as existing techniques only provided limited insights, see Section 3.2. In the end, we bypassed in-depth analysis of FSM state graphs and their interplay with the data path and other FSMs by relying on virtual probing. Obtaining the I/O recordings required for this purpose presented engineering challenges on its own, see Figure 7, but these are situational and depend on the PCB. The main challenge for recovering the implemented algorithm was understanding the dynamic circuit behavior (C6), which almost entirely depended on external data (C5). To this end, simulation results were manually analyzed over many weeks. Finally, understanding the implemented algorithm required domain knowledge beyond our reverse engineering expertise (C7) and required a signal processing expert.
Throughout our case study, validating intermediate results within the black-box setting we were operating in presented another significant challenge. Subsequent stages always depended on previous findings, risking error propagation. This concern was particularly pronounced during word-level reconstruction and algorithmic recovery, where our efforts often felt akin to groping in the dark. Identifying and tracing errors was sometimes impossible. While the SMT-based verification of arithmetic operations offered an initial anchor point, this alone could not fully resolve the issue. In this regard, insights derived from virtual probing and the recorded I/O played a crucial role in achieving end-to-end validation of our reverse engineering result. Virtual probing was also instrumental in interpreting the data flowing through Maggie, enabling us to generate plots of processed data and extract equations representing the implemented algorithm.
An overshadowing issue of our entire case study was the lack of adequate, openly available tooling to automatically investigate the recovered netlist at scale. Developing respective tooling, e.g., for netlist simulation, required immense efforts and was the main reason for many of the months spent on conducting the case study.
RQ2
In light of RQ2, we developed fully automated techniques that address many of the challenges we encountered in our case study. Furthermore, we demonstrated their applicability across FPGA architectures, particularly focusing on Lattice and Xilinx FPGAs. We chose Xilinx as our second evaluation platform as they are the market leader among FPGA vendors and their 7-series FPGAs are more complex than the Lattice iCE40 device family.
Generalizing netlist pre-processing, data-path analysis, and bit-order reconstruction required only minor tweaks, because the techniques used for our case study were already rather independent of the underlying architecture. Here, we merely focused on addressing issues arising from different implementations ranging from CPUs to DSP designs.
While the redesign of the functional arithmetic verification required significant effort and months to complete, adding support for different architectures now is a matter of hours and only requires extending the structural candidate identification. Furthermore, we showed that the labor-intensive dynamic extraction of the algorithm implemented on Maggie from its DSPs could be supported by our guided symbolic execution approach. This significantly reduced the time to get from word-level to an algorithmic description to at most a couple of days.
Still, we noticed differences in the quality of our algorithms’ results based on the analyzed kind of implementation. DSP designs similar to Maggie generally yield better results than the CPUs. We assume that this is because of the more structured data and control path in DSP designs. The higher number of arithmetic operations in a DSP designs compared to a CPU or cryptographic core allows us to automatically identify larger parts of the combinational data path. CPUs leverage varying bits of a register for a wide range of arithmetic operations and thus have a more interwoven data path, which sometimes required us to adapt our techniques. In particular, introducing majority voting to the bit-order propagation massively improved results for CPU implementations.
One challenge, however, remains unsolved: the identification and separation of control logic, see Section 3.2. We demonstrated that this issue can be bypassed using dynamic analysis, but this may not always be feasible, as it requires observing an operational system.
By generalizing our techniques we have demonstrated that some of the netlist reverse engineering challenges can be considered a one-time overhead, because the respective tooling only has to be developed once and can then be applied across FPGA implementations and architectures. Consequently, we contribute to significantly reducing the manual effort of FPGA netlist reverse engineering by providing open-source implementations of the techniques presented in this section, allowing for more a realistic evaluation of countermeasures.
5.2 FPGA Threats & Defenses
Threats of FPGA Reverse Engineering
We showcased the feasibility of reverse engineering a real-world FPGA design, even within the limitations of a resource-constrained academic setting. Compared to regular ICs, the entry bar to reverse engineering is much lower. Unlike FPGAs, for which reliable open-source bitstream documentation projects are readily available [15, 32], the extraction of a netlist from an IC requires complex tooling, expensive equipment, and specialized expertise for preparation, imaging, and image analysis [3, 4, 5, 6, 7]. Given the inherent challenges of IC reverse engineering, we assume that FPGAs are vulnerable to reverse engineering by less powerful adversaries. In contrast, ICs that are more likely to be targeted by sophisticated attackers with nation-state level resources. On top of that, extracting an error-free netlist is much more likely for FPGAs than it is for ICs, since for the latter, error-prone sample preparation and image analysis is needed while an error-free bitstream documentation is more straight-forward to achieve. A correctly recovered gate-level netlist enables techniques like virtual probing (and simulation in general) that are sensitive to errors in the analyzed netlist, thereby increasing the threat potential even further.
In addition to IP theft, large-scale reverse engineering efforts can also provide an entry point for hardware Trojan insertion given the reconfigurability of FPGAs [48, 49, 50]. Consequently, a natural defense against netlist extraction would be to not use FPGAs for sensitive IP altogether. We recognize that this is not a viable option in many cases, hence additional FPGA countermeasures should be considered to defend against IP theft and Trojan insertion.
Bitstream Protections
A first line of defense is to impede extraction of a gate-level netlist. Proprietary bitstream formats may initially raise the bar for a reverse engineering attack, but given increasingly automated efforts, this may be overcome within reasonable time. This is inevitable since the user will always be able to translate HDL designs into bitstreams using EDA tools. As offered by major FPGA vendors, reliable cryptographic schemes can ensure bitstream confidentiality, integrity, and authentication. By encrypting and signing a bitstream, attacks such as IP theft are impeded. However, such bitstream protections are not always provided on low-cost FPGAs and legacy devices, which may continue to be in use for decades.
Netlist Protections
In light of recent attacks on bitstream encryption schemes [21, 16, 17, 18, 19, 20], further precautions must be taken to harden the netlist against static and dynamic analysis by deteriorating those netlist properties that facilitate automated reverse engineering. Effective netlist obfuscation must aim at preventing the recovery of word-level structures as these structures are crucial for further analysis [51]. Dynamic reverse engineering is key to achieve an algorithmic understanding of larger circuits. Hence, to harden against such analysis techniques, non-simulatable primitives such as physical unclonable functions [52] or partial reconfiguration [38, 53, 38] should be included in the design and merged into the control and data paths.
5.3 Related Work
Existing netlist reverse engineering techniques often only tackle isolated problems and have been evaluated on (sometimes outdated) open-source benchmarks rather than black-box designs. Hence, a comprehensive real-world case study is still lacking from literature. For a summary of prior work on netlist reverse engineering, see Azriel et al. [23].
Revisiting Related Work
One research strand focuses on automated state registers recognition and FSM recovery [39, 35, 54, 55, 34]. RELIC by Meade et al. [37, 34] rates the similarity of FF fan-in trees to determine state and non-state FFs. Chowdhury et al. [56] propose a graph neural network (GNN) base approach that can also separate control FFs from data. To identify known sub-circuits, functional [57, 58, 59, 34] and graph-similarity based approaches [60] have been proposed. WordRev by Li et al. [58] comprises methods to recover word-level structures for functional matching. Subramanyan et al. [59] builds on their work to identify components such as adders, multipliers, counters, and registers. Albartus et al. introduced DANA [42] to identify high-level registers in a gate-level netlist. Meade et al. presented REBUS and REWIND to recover the datapath using similarity metrics comparable to RELIC [34]. Alrahis et al. [61] utilize GNN to determine for each gate whether it belonged to a module that implements known functionality. Narayanan et al. [47] also propose an identification of arithmetic operations by taking advantage of carry chains on FPGAs, as discussed in Section 4.2. Some work has been done to partition a netlist based on the inter connectivity inside a cluster of gates. Werner et al. [62] evaluate the applicability of conventional graph clustering algorithms for clustering in circuits, while Hong et al. [63] search for partitions by optimizing a n-cut with a GNN.
Advocating for Open-Source Implementations
In the software community, there is an agreement on common (open-source) tools and methods. This is not the case for hardware reverse engineering. Existing netlist reverse engineering techniques often lack reference implementations, posing challenges to reproducibility. During our case study, we encountered difficulties in applying existing techniques due to the effort required for re-implementation, surpassing the scope of our work.
We acknowledge the complexity of this issue, understanding that providing open-source implementations may clash with industry and government interests, occasionally even triggering concerns related to export regulations. We have personally encountered similar challenges. Despite these considerations, we advocate for a more open approach to contribute to the collective progress in netlist reverse engineering research, enabling the FPGA community to tackle more complex case studies, such as entire SoCs, based on our insights. This transparency is crucial for fostering a comprehensive understanding of the associated threats. We assume that such efforts are mostly conducted by governments behind closed doors, often away from public scrutiny.
5.4 Limitations and Future Work
Our case study, while focused on a single DSP design on a relatively small FPGA, still provides valuable insights. As we have shown in Section 4, our case study findings inform generalized techniques that are applicable to implementations across different architectures, designs, and sizes. Future research could build up on this work to explore more elaborate real-world designs on more complex FPGAs to further underline the threat potential of IP theft for FPGAs. Our work builds the foundation for such case studies in that it provides the algorithms and implementations to dissect such a netlist at scale. Also, we have shown that there still is a need for better automated techniques for control logic extraction and analysis in gate-level netlists.
6 Conclusion
Our work demonstrates the challenges associated with IP theft through FPGA reverse engineering and discusses the challenges faced in a real-world setting. To this end, our case study on a Lattice iCE40 FPGA that is part of iPhone 7 underscores the vulnerability of FPGAs to IP theft due to the low barrier for netlist extraction compared to regular ICs. Thereby, we reveal that while IP theft from FPGAs requires substantial effort, specialized skills, and domain knowledge, some of the most labor-intensive steps can be automated. We contribute to the field by introducing generalized netlist reverse engineering techniques and open-source implementations, which reduce manual effort and facilitate future research in this domain. Our techniques prove effective across different FPGA architectures and highlight the threat of IP theft even by less resourceful adversaries. This work emphasizes the importance of robust bitstream and netlist protections to safeguard valuable IP in FPGAs. Finally, we call for more open and transparent research in FPGA reverse engineering to better understand and mitigate associated threats.
Acknowledgments
We thank Marc Fyrbiak and Max Hoffmann for the fruitful discussions. Funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany´s Excellence Strategy - EXC 2092 CASA - 390781972, through ERC grant 695022, NSF grants CNS-1563829 and CNS-1749845, and ISF project IZ25-5793-2019-43 (“CHIoSec”).
Appendix A Simulation in HAL
For netlist simulation in HAL, e.g., in the context of virtual probing, we extended the framework with an interface for commercial-grade simulation tools such as Verilator [64]. Furthermore, we added a waveform viewer to the HAL graphical user interface (GUI) that can be controlled via a Python application programming interface (API), see Figure 13. The dynamic state of the netlist can also be represented by coloring the nets in the graph view according to their current signal value, which adapts whenever the cursor in the waveform viewer is moved. See Figure 11 and Figure 12 for pictures of our actual virtual probing setup as initially described in Figure 7.
Figure 11: Side view of our custom breakout board sitting between the iPhone PCB and Maggie.
Figure 12: Top view of a Saleae logic analyzer being connected to the iPhone via our custom breakout board.
Figure 13: Screenshot of HAL showing the module tree on the left, the graph view in the middle, and the simulator on the right. The nets are colored depending on their current value, blue represents a 0 and red a 1.
Appendix B Details on the Maggie DSP
In the following, we briefly comment on three key building blocks of the DSP implementation, as shown in Figure 9: $(i)$ an input processing, $(ii)$ a state observer to track the actuator dynamics, and $(iii)$ a closed-loop controller making the actuator follow a desired movement.
Input Processing
Maggie’s data path takes two input signals supplied via Inter-IC Sound (I2S) and Serial Peripheral Interface (SPI) from the motor driver and Homer, respectively. These two signals are combined (scaled and subtracted) and then passed to a look-up table-based function where linear interpolation is used to enhance resolution. We believe that this step translates sensor measurements to the current LRA position, relying on calibration.
State Observer
The input processing result and the closed-loop controller output are passed to a processing stage which implements a state observer, estimating the LRA’s position and velocity. In particular, Maggie appears to implement a proportional-integral (PI) observer [65].
Closed-Loop Controller
The closed-loop controller takes a time series of desired actuator positions as its reference signal. Thus, as different haptic events are to be generated, this signal is changed. For instance, it is updated with the haptic feedback setting visible in Figure 12. Using the reference signal and the state observer outputs, the closed-loop controller accumulates the position and the velocity error (the desired velocity is obtained from numerical differentiation of the reference signal) and subtracts the integrated position estimation error. The particular combination of error components suggests that the control loop is a proportional-integral-derivative (PID)-controller [66, 67, 68]. Before being passed to the motor driver via I2S, a low-pass filter is applied to the output of the closed-loop controller.
Appendix C MUX Evaluation
For the evaluation of our word-level MUX detection using DANA, refer to Figure 14. The figure illustrates the five most frequently recovered MUX sizes for each benchmark, once synthesized for a Xilinx 7-series FPGA and once for s Lattice iCE40 FPGA. In this analysis, we exclusively consider MUXes comprising at least four gates. For the hilbert_transformer, no word-level MUXes were identified, hence we omitted it in Figure 14. We also provide our results for Maggie, which are only available for Lattice.
(a) ibex (Expected: 32)
(b) canny_edge (Expected: 8,16,32)
(c) icicle (Expected: 32)
(d) simple_risc_v (Expected: 32)
(e) fft64 (Expected: 19,18,17,16)
(f) sha256 (Expected: 256,32)
(g) Maggie (Expected: 16)
Figure 14: Evaluation of our word-level MUX detection using DANA.